OpenStudy (anonymous):

A tuning fork is sounded above several pipes closed at one end. Two pipes with lengths of 20 cm and 60 cm produce the 1st and 2nd resonances respectively. What is the frequency of the fork if the speed of sound in air is 340 m/s?

OpenStudy (anonymous):

For open ended pipes resonance hits at lenghts w/4 , 3w/4 , 5w/4 where w is the wavelenght of the sound produced. Now for ist resonance L=w/4 . Here L=20. Find w and then use relation frequency , v= V/w where V is velocity.

OpenStudy (anonymous):

oh... thank you so much! ^^